# Large Language Models Are Cross-Lingual Knowledge-Free Reasoners

### v1

@@ 4.4.1 Overall computational similarity
In this section, we assess the LLaMA-2-7B-Chat model’s hidden states cosine similarity and Feed-Forward Network neuron activation overlap on knowledge-related tasks and knowledge-free reasoning tasks. We also evaluate changes in these metrics after fine-tuning with LoRA to understand the mechanisms supporting cross-lingual transfer of knowledge-free reasoning abilities. The experimental results are shown in Figures 7 and 8.
Figure 6: Averaged XLTR from English to Arabic/Hebrew on KFRD for models in different stages trained in Arabic/Hebrew
@@
Figure 7: CS for different datasets in the LLaMA-2-7B-Chat model. The black lines on each bar indicate the 99% confidence intervals estimated with bootstrap sampling (Efron, 1992).
@@
Figure 8: NAO for different dataset in the LLaMA-2-7B-Chat at activation thresholds ranging from 0.1 to 0.9. Shaded areas: 99% confidence intervals estimated with bootstrap sampling; Solid lines: results of the original model; Dashed lines: results of the LoRA tuned model. The meanings of the shaded areas and dashed lines in Figures 9 and 10 are consistent with those described here.
@@
Internal representation of knowledge-free reasoning task is better aligned than retrieval
@@
The results in Figure 7 indicate that the CS of the model on knowledge-free reasoning tasks is significantly higher than that on knowledge-related reasoning tasks both before and after training. Additionally, after training on the knowledge-free reasoning dataset, the model’s CS shows a significant increase. In contrast, after training on the knowledge-related reasoning dataset, the increase in CS is not significant and may even decrease. This suggests that adapting to knowledge-free reasoning tasks results in a more aligned hidden space processing across different languages, whereas adapting to knowledge-related reasoning tasks leads to relatively divergent hidden space processing.
@@
Neuron activation pattern of knowledge-free reasoning task is more similar than retrieval
@@
Neuron analysis further elucidates this phenomenon. The results in Figure 8 show that, across all threshold settings, NAO for knowledge-free reasoning tasks is significantly higher than for knowledge-related reasoning tasks. This indicates that the model tends to use similar neurons for processing knowledge-free reasoning tasks across different languages, resulting in similar neuron activation patterns and computational processes. Consistent with the hidden states results, after training on the knowledge-free reasoning dataset, the model’s NAO shows a significant increase, whereas there is no significant improvement and even a decline after training on the knowledge-related reasoning dataset. This suggests that training on knowledge-free reasoning tasks makes neuron activation characteristics across different languages more similar, whereas training on knowledge-related reasoning tasks makes these characteristics relatively divergent.
@@
These results provide a comprehensive analysis of the differing cross-lingual transfer effectiveness between knowledge-free reasoning and knowledge-related reasoning from a computational similarity perspective. We hypothesize that this difference is because the model tends to store knowledge for different languages in different neurons, while using similar neuron groups for knowledge-free reasoning functions.
@@ 4.4.2 Layer-wise computational similarity
To analyze the relationship between hidden states and neuron activation characteristics with cross-lingual transfer of reasoning abilities more granularly, we performed a layer-wise analysis of the cross-lingual cosine similarity of hidden states and cross-lingual neural activation overlap. The experimental results are shown in Figures 9 and 10.
@@
It is observed that the phenomena identified in the previous section, specifically the significantly higher CS and NAO for knowledge-free reasoning tasks compared to knowledge tasks with more pronounced improvements after fine-tuning, are most evident in the middle layers (layers 6-25) of the model and less in the shallow layers (layers 1-6). Previous work (Zhao et al., 2024; Wendler et al., 2024) suggested that the middle layers of LLMs are primarily responsible for conceptual reasoning, which is cross-lingual. This hypothesis aligns with our findings and further supports the view that reasoning capabilities can transfer across languages.
Additionally, in the upper layers (layers 25-32) of the model, the CS and NAO for knowledge-free reasoning tasks are similar to those for knowledge tasks before training, but show significant improvement after training. This phenomenon is accompanied by a consistent performance improvement across different languages. The layered functionality theory of LLMs suggests that the upper layers are mainly responsible for generating the tokens of the target language responses. Therefore, we infer that task fine-tuning makes the model’s behavior more consistent during the decoding phase, resulting in better multilingual task performance.
@@ 5 Related Work
Laskar et al. (2023) performed evaluation for multilingual ability of ChatGPT. Shi et al. (2022) found LLMs can perform reasoning in multiple languages using CoT, even for those languages with very low resources.
@@
Gao et al. (2024) evaluated the cross-lingual transferability of models on multiple reasoning datasets, finding significant variations in transfer performance across different datasets. Additionally, Zhu et al. (2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks. Ye et al. (2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge. Furthermore, Hu et al. (2024) found that knowledge transferability remains weak across various settings.
@@
Zhao et al. (2024) analyzed the way LLMs handle multilingualism and suggested a three-phase working pattern, which includes understanding, task solving and generation. Wendler et al. (2024) also arrived at a similar conclusion.
6 Conclusion
In this study, we analyze the reasons behind the differing cross-lingual transfer abilities of LLMs on various reasoning tasks. We divide reasoning tasks into two components: knowledge retrieval and knowledge-free reasoning. We adapt several existing reasoning datasets and construct a knowledge-free reasoning dataset for analysis.
Experiments on these datasets demonstrated that the requirement of knowledge retrieval significantly lowers the cross-lingual transfer performance of reasoning tasks, while the knowledge-free reasoning ability can nearly fully transferred between languages.
In addition, we analyze the difference of computational processes between knowledge-related and knowledge-free reasoning tasks to explain the gap in cross-lingual transfer. The results show that the overall cross-lingual computational similarity of knowledge-free reasoning is significantly higher than that of knowledge-related reasoning, particularly in the middle-high layers of the model. This suggests that knowledge-free reasoning benefits from a shared neural mechanism across languages, whereas knowledge storage is more language-specific.
@@ 6 Conclusion
Appendix A Detailed Description of Knowledge-Free Reasoning
The KFRD is generated using a unified template, consisting entirely of multiple-choice questions with four options. Each question is structured into three parts: input, output, and transformation rules. Specific examples can be seen in Table 1, and the templates used for these examples are shown in Figure 11.
Arithmetic Reasoning
Input 11,645 (one or two numbers)
Transformation Rule Addition (a mathematical operation)
Output Options A) 595 B) 536 C) 771 D) 656 (one number, sometimes the result is two numbers)
Symbolic Reasoning
Transformation Rule Swap the positions of the 5th and 2nd words; Delete the 2nd word (symbolic operations within three layers)
Logical Reasoning
Input Alex is Aurora Vale. Everything that is Aurora Vale is Omicron Delta. Stella is not Chronos Wasteland. Max is not Dreamweaver’s Haven. Suppose Sally is Whispering Meadows, then Sally is Chimerical Citadel. Everything that is Ebonwyrm Abyss is Phoenixfire Ridge. (8 propositions)
Transformation Rule Implication Elimination (a logical rule)
Table 1: Examples of different tasks in the KFRD dataset
A.1 Arithmetic Reasoning
This dataset transforms two input numbers through mathematical operations into one or two output numbers. The mathematical operations include basic arithmetic operations (addition, subtraction, multiplication, division), equality, geometric progression, arithmetic progression, and sorting.
@@ A.1 Arithmetic Reasoning
A.2 Symbolic Reasoning
This dataset transforms 3-5 input words from the corresponding language through symbolic operations to generate the output. Symbolic operations include repetition, addition, deletion, reordering, and their combinations. Considering that single-step symbolic operations are too simple, we chose up to three-step symbolic operations.
@@ A.2 Symbolic Reasoning
A.3 Logical Reasoning
This dataset generates output from a subset of eight input propositions using logical rules. Logical rules include Implication Elimination, Conjunction Introduction, Conjunction Elimination, Disjunction Introduction, Disjunction Elimination, and Proof by Contradiction. The Logical rules are referenced from Saparov et al. (2024).
@@ A.3 Logical Reasoning
Figure 11: Example prompt template for our KFRD dataset
@@
Table 2: Training models for Arabic and Hebrew
Appendix B Dataset Examples
We provide some examples of adapted datasets used in this paper in Table 3.
StrategyQA
Answer True
MKQA
Query Who sings "I Hear You Knocking But You Can’t Come In"?
Answers Dave Edmunds
BoolQ
Question Do Iran and Afghanistan speak the same language?
Answer True
AmbigQA
Answer 74 days
Table 3: Examples of adapted datasets used in this paper
Appendix C Language Choice
This section provides an overview of the languages utilized in our research, highlighting the primary countries where they are spoken and their respective language families. Refer to Table 4 for detailed information.
ISO Countries Language Family
@@ Appendix C Language Choice
Table 4: Correspondence between Languages, Countries, and Language Families
Appendix D Implementation Details for Interpretability
D.1 Dataset adjustments
We made slight modifications of the datasets to ensure that the last character of different datasets is uniformly a "?" token. Since we are analyzing the internal representation of the last token, in this way, we can eliminate interference caused by the inconsistent input token, which may make the representation unreliable, especially in the bottom layers. Another reason why we append the token "?" is that it can act as a trigger to let the model start the process of preparing to answer the question, which is what we are analyzing.
@@ D.2 Calculation method for activation values
Table 5: Training epoch and number of samples of fine-tuned datasets in the transferability experiments
@@
Table 6: The size of testset used in the transferability experiments
Appendix E Experiments Details
@@ E.2 Hyperparameters
The numbers of training epoch and samples used in the transferability experiments are listed in Table 5. These numbers are tuned to enable LLaMA-2-7B-Chat to achieve 85% + accuracy on the corresponding tasks. The size of testsets used in the transferability experiments are shown in Table 6.
@@
E.3 Computation resources
@@
Appendix F Models Used in the Target Language Proficiency Experiment
@@
The continue pre-training or fine-tuning models of LLaMA-2-7B and Mistral-7B used in the target language proficiency experiment in 4.3.2 are listed in Table 2.
Appendix G Additional Results of Experiment
Here we provide the accuracy of the above experiments in Figure 12, 13 and 14.

### v2

@@ 5.4.1 Overall computational similarity
In this section, we assess the original and fine-tuned LLaMA-2-7B-Chat model’s CS and NAO on knowledge retrieval and knowledge-free reasoning tasks. The experimental results are shown in Figures 8 and 9.
@@
Figure 8: CS for different datasets in the LLaMA-2-7B-Chat model. Black lines on each bar indicate the 99% confidence intervals estimated with bootstrap sampling (Efron, 1992).
@@
Figure 9: NAO for different dataset in the LLaMA-2-7B-Chat at activation thresholds ranging from 0.1 to 0.9. Shaded areas: 99% confidence intervals estimated with bootstrap sampling; Solid lines: results of the original model; Dashed lines: results of the LoRA tuned model. The meanings of the shaded areas and dashed lines in Figures 10 and 11 are consistent with those described here.
@@
Internal representation of knowledge-free reasoning task is better aligned than knowledge retrieval
@@
The results in Figure 8 indicate that the CS of the model on knowledge-free reasoning tasks is significantly higher than that on knowledge retrieval tasks both before and after fine-tuning. Additionally, after fine-tuning on knowledge-free reasoning datasets, the CS increases significantly on the corresponding datasets, while fine-tuning on knowledge retrieval datasets shows no significant improvement and may even lead to a decrease. This suggests that adapting to knowledge-free reasoning tasks results in a more aligned hidden space processing across languages.
@@
Neuron activation pattern of knowledge-free reasoning task is more similar than knowledge retrieval
@@
Neuron analysis further elucidates this phenomenon. The results in Figure 9 show that, across all activation threshold settings, NAO for knowledge-free reasoning tasks is significantly higher than for knowledge retrieval tasks. This indicates that the model tends to use similar neurons for processing knowledge-free reasoning tasks across different languages, resulting in similar neuron activation patterns. Consistent with the hidden states results, after training on the knowledge-free reasoning dataset, NAO increases significantly, whereas there is no significant improvement and even a decline after training on the knowledge retrieval dataset. This suggests that training on knowledge-free reasoning tasks makes neuron activation characteristics across different languages more similar, leading to the conclusion that the knowledge-free reasoning ability share a similar set of neurons.
@@
These results provide a comprehensive analysis of the different cross-lingual transfer effectiveness between knowledge-free reasoning and knowledge retrieval component from a computational similarity perspective. We hypothesize that this difference is because the model stores knowledge for different languages in different neurons, while using similar neuron groups for knowledge-free reasoning.
@@ 5.4.2 Layer-wise computational similarity
To gain finer-grained insights, we performed a layer-wise analysis of CS and NAO. The experimental results are shown in Figures 10 and 11.
@@
It is observed that the significantly higher CS and NAO for knowledge-free reasoning tasks, compared to knowledge retrieval tasks, are most pronounced in the middle layers (layers 6-25). Previous work (Zhao et al., 2024; Wendler et al., 2024) suggested that the middle layers of LLMs are primarily responsible for conceptual reasoning, which is cross-lingual. This hypothesis aligns with our findings and further supports the view that knowledge-free reasoning capabilities can transfer across languages.
Additionally, the upper layers (26-32) show similar CS and NAO patterns for both knowledge-free reasoning and knowledge retrieval tasks before training, but training improvements are only notable in knowledge-free reasoning. We find that fine-tuning on knowledge-free tasks significantly enhances multilingual accuracy, leading to more consistent outputs. Since the upper layers primarily handle token generation (Zhao et al., 2024; Wendler et al., 2024), this consistency improvement results in higher CS and NAO.
@@ 6 Related Work
Laskar et al. (2023) performed evaluation for multilingual ability of ChatGPT. Shi et al. (2022) found LLMs can perform reasoning in multiple languages using CoT, even for those languages with very low resources. Their analysis mainly evaluated different reasoning tasks, but did not investigate the reasons for performance variations.
@@
Gao et al. (2024) evaluated the cross-lingual transferability of models on multiple reasoning datasets, finding significant variations in transfer performance across different datasets. Furthermore, Hu et al. (2024) found that knowledge transferability remains weak across various settings. Building on their conclusions, we distinguish between the knowledge retrieval and knowledge-free reasoning components and extend the analysis to all reasoning tasks. We include further discussion of related reasoning works in Appendix H.
@@
Zhao et al. (2024) analyzed the way LLMs handle multilingualism and suggested a three-phase working pattern, which includes understanding, task solving and generation. Wendler et al. (2024) also arrived at a similar conclusion. Expanding on their findings, we further analyzed the differences in how LLMs handle reasoning and knowledge tasks across languages.
7 Conclusion and Discussion
In this study, we analyze the reasons behind the differing cross-lingual transfer abilities of LLMs on various reasoning tasks. We divide reasoning tasks into two components: knowledge retrieval and knowledge-free reasoning. Our experiments demonstrated that the demand for knowledge retrieval significantly hinders the cross-lingual transfer performance, while the knowledge-free reasoning ability can be nearly fully transferred between languages. This discrepancy arises because knowledge-free reasoning relies on shared neural mechanisms across languages, while knowledge storage tends to be more language-specific.
Based on these findings, for knowledge, we recommend prioritizing the inclusion of multilingual data in training corpora in the future. For reasoning, emphasis should be placed on the quality of reasoning data rather than the number of languages. Furthermore, for future multilingual analysis, we recommend investigating knowledge retrieval and knowledge-free reasoning components individually to gain more targeted and detailed insights.
@@ 7 Conclusion and Discussion
Appendix A Detailed of Dataset
A.1 Detailed description of Knowledge-Free Reasoning Dataset
The KFRD is generated using a unified template, consisting entirely of multi-choice questions with four options. We first create parallel templates for 10 languages using GPT-4 and then fill in different parts of the template with pre-defined rules. Each question is structured into three parts: input, output, and transformation rules. Specific examples can be seen in Table 1, and the templates used for these examples are shown in Figure A1.
A.1.1 Arithmetic reasoning
This dataset transforms two input numbers through mathematical operations into one or two output numbers. The mathematical operations include addition, subtraction, multiplication, division, equality, geometric progression, arithmetic progression, and sorting. Each of the three parts are generated by the following rules:
@@ A.1.1 Arithmetic reasoning
A.1.2 Symbolic reasoning
This dataset transforms 3-5 input words from the corresponding language through symbolic operations to generate the output. Symbolic operations include repetition, addition, deletion, reordering, and their combinations. Considering that single-step symbolic operations are too simple, we chose up to three-step symbolic operations. Each of the three parts are generated by the following rules:
@@ A.1.2 Symbolic reasoning
A.1.3 Logical reasoning
This dataset generates output from a subset of eight input propositions using logical rules. Logical rules include Implication Elimination, Conjunction Introduction, Conjunction Elimination, Disjunction Introduction, Disjunction Elimination, and Proof by Contradiction. The Logical rules are referenced from Saparov et al. (2024). Each of the three parts are generated by the following rules:
@@ A.1.3 Logical reasoning
Figure A1: Example prompt template for our KFRD dataset
@@
Table A1: Training models for Arabic and Hebrew
A.2 Detail of existing pseudo knowledge-free reasoning datasets
Here we provide more details on the datasets used in the experiment.
A.3 Translation process for English-only datasets
For datasets available only in English, we translate them into other languages with Google Translate and verify translation quality with GPT-4.
Google Translate is highly regarded in the field of commercial translation and is widely used in multilingual research (Chen et al., 2021; Ye et al., 2023; Omar and Salih, 2024; Song et al., 2024). To ensure translation accuracy, we sampled a subset of translation results and employed GPT-4 for verification. Due to budget constraints, we were unable to employ human translators.
For the StrategyQA dataset, we utilized Google Translate and conducted a sample check of 100 items for each language using GPT-4. This process resulted in an overall quality score of 4.47 (on a scale of 1-5), which we consider acceptable for our purposes.
Appendix B Language Choice
This section provides an overview of the languages utilized in our research, highlighting the primary countries where they are spoken and their respective language families. Refer to Table A2 for detailed information.
ISO Country Samples Language Family
@@ Appendix B Language Choice
Table A2: Correspondence between Languages, Country Samples, and Language Families
Appendix C Implementation Details for Interpretability
C.1 Calculation method for activation values
C.2 Reasons for using the last token for analysis
C.3 Dataset adjustments
To ensure consistency in the final token across different datasets, we made slight modifications by adding a language-specific "?" where needed.
Since we are analyzing the internal representation of the last token, in this way, we can eliminate interference caused by the inconsistent input token, which may make the representation unreliable, especially in the bottom layers. Another reason why we append the token "?" is that it can act as a trigger to let the model start the process of preparing to answer the question, which is what we are analyzing.
@@ C.3 Dataset adjustments
Table A3: Training epoch and number of samples of fine-tuned datasets in the transferability experiments
@@
Table A4: The size of testset used in the transferability experiments
Appendix D Experiments Details
@@ D.2 Hyperparameters
The numbers of training epoch and samples used in the transferability experiments are listed in Table A3. These numbers are tuned to enable LLaMA-2-7B-Chat to achieve 85% + accuracy on the corresponding tasks. The size of testsets used in the transferability experiments are shown in Table A4.
@@
StrategyQA
Answer Yes
QASC
Question Climate is generally described in terms of what?
Facts 1. Climate is generally described in terms of temperature and moisture. 2. Fire behavior is driven by local weather conditions such as winds, temperature and       moisture.
Options A. sand B. occurs over a wide range C. forests D. Global warming E. rapid changes occur F. local weather conditions G. measure of motion H. city life
Answer F
Table A5: Examples of knowledge-involved datasets
D.3 Computation resources
@@
D.4 Models used in the target language proficiency experiment
@@
The continue pre-training or fine-tuning models of LLaMA-2-7B and Mistral-7B used in the target language proficiency experiment in 4.3.2 are listed in Table A1.
Appendix E Additional Results of Experiment
Here we provide the accuracy of the above experiments in Figure A2, A5, A6 and A7.
We provide the results of the QASC dataset in Figure A3 and A4. The results show that the more knowledge provided leads to better cross-lingual transferability, which aligns with our conclusion.
Appendix F Language Distribution of Model Training Corpora
In this section, we present the language distribution of the pre-training corpora, referencing Table A6 from the LLaMA2 paper (Touvron et al., 2023) and Table A7 from the BLOOM paper (Workshop et al., 2022). Unfortunately, we were unable to locate the corresponding distribution data for Mistral and Qwen.
For LLaMA2, languages such as Arabic and Hebrew were not included in the table, indicating that their proportions are lower than 0.005%, categorizing them as extremely low-resource languages. The other eight languages discussed in the paper are represented. Notably, German and Chinese rank as high-resource languages, accounting for 0.17% and 0.13% of the corpus, respectively, holding the second and fifth highest positions.
For BLOOM, only English, French, Chinese, and Arabic are explicitly listed, while other languages are not reported in the table.
Language Percent Language Percent
en 89.70% uk 0.07%
unknown 8.38% ko 0.06%
de 0.17% ca 0.04%
fr 0.16% sr 0.04%
sv 0.15% id 0.03%
zh 0.13% cs 0.03%
es 0.13% fi 0.03%
ru 0.13% hu 0.03%
nl 0.12% no 0.03%
it 0.11% ro 0.03%
ja 0.10% bg 0.02%
pl 0.09% da 0.02%
pt 0.09% sl 0.01%
vi 0.08% hr 0.01%
Table A6: Language distribution in pretraining data with percentage >= 0.005%. Most data is in English, meaning that Llama 2 will perform best for English-language use cases. The large unknown category is partially made up of programming code data.
Language Size in Bytes Language Size in Bytes
Akan 70,1554 Arabic 74,854,900,600
Assamese 291,522,098 Bambara 391,747
Basque 2,360,470,848 Bengali 18,606,823,104
Catalan 17,792,493,289 Chichewa 1,187,405
chiShona 6,638,639 Chitumbuka 170,360
English 484,953,009,124 Fon 2,478,546
French 208,242,620,434 Gujarati 1,199,986,460
Hindi 24,622,119,985 Igbo 14,078,521
Indonesian 19,972,325,222 isiXhosa 14,304,074
isiZulu 8,511,561 Kannada 2,098,453,560
Kikuyu 359,615 Kinyarwanda 40,428,299
Kirundi 3,272,550 Lingala 1,650,804
Luganda 4,568,367 Malayalam 3,662,571,498
Marathi 1,775,483,122 Nepali 2,551,307,393
Northern Sotho 1,764,506 Odia 1,157,100,133
Portuguese 79,277,543,375 Punjabi 1,572,109,752
Sesotho 751,034 Setswana 1,502,200
Simplified Chinese 261,019,433,892 Spanish 175,098,365,045
Swahili 236,482,543 Tamil 7,989,206,220
Telugu 29,934,071,59 Traditional Chinese 762,489,150
Twi 1,265,041 Urdu 2,781,329,959
Vietnamese 43,709,279,959 Wolof 3,606,973
Xitsonga 707,634 Yoruba 89,695,835
Programming Languages 174,700,245,772
Table A7: Linguistic makeup of the ROOTS corpus. The ROOTS corpus is a multilingual dataset used for pretraining the BLOOM model.
Appendix G Reasons for Creating a New Dataset
The primary reason for creating a new dataset is that most existing datasets involve knowledge retrieval, which does not align with our focus on knowledge-free reasoning. For instance, in StrategyQA, while necessary reasoning knowledge is provided, it may be incomplete.
StrategyQA Example:
  • Question: Are you likely to find a crucifix in Karachi?
  • Facts: The crucifix is a symbol of Christianity. The vast majority of Pakistan’s population is Muslim.
  • Missing Knowledge: It is not specified that Karachi is in Pakistan.
Similarly, most existing math datasets also require knowledge retrieval to answer questions, such as the ASDiv-a dataset.
ASDiv-a Example 1:
  • Question: At the school’s book fair, Sam bought 13 adventure books and 17 mystery books. If 15 of the books were used, how many new books did he buy?
  • Missing Knowledge: The new books are those that were not used.
ASDiv-a Example 2:
  • EN-Question: After the aviary was the zoo’s swamp area. Penny counted a total of 55 tree frogs, 10 poison frogs, and 13 wood frogs. How many frogs was Penny able to count?
  • FR-Question: Après la volière se trouvait la zone marécageuse du zoo. Penny a dénombré un total de 55 rainettes, 10 grenouilles venimeuses et 13 grenouilles des bois. Combien de grenouilles Penny était-elle capable de compter ?
  • Missing Knowledge: In English, it can be inferred that “poison frogs," “wood frogs," and “tree frogs" are all “frogs." However, in French, it is not directly inferable that “rainettes" are a type of “grenouilles," requiring additional knowledge retrieval.
Some existing logic datasets are not designed with knowledge-free reasoning in mind, as they use real-world entities. This leads to situations where, although it is theoretically possible to answer without retrieving external knowledge, the retrieval of such knowledge might influence the final answer. For example, given the statement “Harry is a cat," the model might infer “Harry is an animal" based on its existing knowledge, without requiring contextual reasoning rules. Similarly, based on prior knowledge, the model might incorrectly assume “The squirrel likes the squirrel" as related, especially when the actual context is irrelevant.
This issue becomes more pronounced when translation is involved. For instance, when translating from English to Chinese, “The squirrel likes the squirrel" may become “squirrels like squirrels," as Chinese does not use articles. This can amplify the influence of pre-existing knowledge, leading to incorrect answers.
By constructing our own dataset, we also avoid potential translation issues that arise when existing datasets are used in different languages, ensuring that reasoning tasks are uniformly understood across languages.
Another advantage of creating a new dataset is that we can control the difficulty level. If the dataset is too difficult and models have low accuracy in English, it would be meaningless to measure cross-lingual transferability. Moreover, a new dataset allows for a more comprehensive coverage of reasoning operations.